# Inspection and Control of Self-Generated-Text Recognition Ability in Llama3-8b-Instruct

### v1

@@ Abstract
It has been reported that LLMs can recognize their own writing. As this has potential implications for AI safety, yet is relatively understudied, we investigate the phenomenon, seeking to establish whether it robustly occurs at the behavioral level, how the observed behavior is achieved, and whether it can be controlled. First, we find that the Llama3-8b–Instruct chat model - but not the base Llama3-8b model - can reliably distinguish its own outputs from those of humans, and present evidence that the chat model is likely using its experience with its own outputs, acquired during post-training, to succeed at the writing recognition task. Second, we identify a vector in the residual stream of the model that is differentially activated when the model makes a correct self-written-text recognition judgment, show that the vector activates in response to information relevant to self-authorship, present evidence that the vector is related to the concept of “self” in the model, and demonstrate that the vector is causally related to the model’s ability to perceive and assert self-authorship. Finally, we show that the vector can be used to control both the model’s behavior and its perception, steering the model to claim or disclaim authorship by applying the vector to the model’s output as it generates it, and steering the model to believe or disbelieve it wrote arbitrary texts by applying the vector to them as the model reads them.
@@ 1 Introduction
Panickssery et al. (2024), utilizing a summary writing/recognition task, report that a number of LLMs, including Llama2-7b-chat, show out-of-the-box (without fine-tuning) self recognition abilities. However, that work focussed on the relationship between self-recognition task success and self-preference, rather than the specific means by which the model was succeeding at the task. Laine et al. (2024), as part of a larger effort to provide a foundation for studying situational awareness in LLMs, utilized a more challenging text continuation writing/recognition task and demonstrate self-recognition abilities in several larger models (although not Llama2-7b-chat), but there the focus was on how task success could be elicited with different prompts and in different models. Thus we seek to fill a gap in understanding what exactly models are doing when they succeed at a self recognition task.
We first demonstrate Llama3-8b–Instruct self-recognition task success in a variety of domains. We are particularly interested in distinguishing “true” self recognition of writing - entailing knowledge of one’s own writing style - from “discriminability” - being able to detect consistent differences in the styles of two sets of texts, or being able to identify texts as being more or less “human-like” (i.e., like the training data). To understand whether the model is engaging in “true” self recognition, which would have the implications for AI safety described above, we next attempt to eliminate competing hypotheses.
Having done so, to help understand, and potentially control, the model representations underlying this self-recognition ability, we apply the contrastive pairs method (Turner et al. (2024), Zou et al. (2023)) to isolate vectors in the residual stream of the model that are distinctively activated for self- vs human-written texts in the context of a paradigm that prompts the model to make a binary judgement about its authorship of a given text. Via inspection with the Tuned Lens Belrose et al. (2023) and a series of steering experiments, we then identify one particular vector that appears to be strongly related to the model’s ability to correctly claim or deny authorship. After further work to characterize the information it carries, we then demonstrate that the vector can be applied to the output token to cause the model to assert or deny authorship at essentially 100% rates for new and out-of-distribution texts in this paradigm, and that the model is much less likely to assert authorship when the vector is completely projected out of the residual stream. Having established that the vector directly affects behavior, we then probe whether it can affect perception, by adding it to or subtracting it from the texts being evaluated, and not doing anything to the output token, in both this paradigm and one in which the model is asked to choose which of two texts it wrote. We find a strong effect of this perceptual “coloring” on the model’s behavior.
@@ 2.1 Testing model self-recognition ability
To generate writing samples we employ two different tasks, referred to hereafter as “Summarization” and “Continuation”. In the Summarization task, following Panickssery et al. (2024), models are prompted to generate summaries or other descriptive information from articles, and are subsequently tested on their ability to distinguish those from responses generated by humans or other models. In the Continuation task, following Laine et al. (2024) and using a portion of the dataset they constructed (“SAD”), we show the models the beginning of a piece of human-generated text, then prompt them to continue it in a plausible manner, then probe their ability to distinguish what they generated from the true continuation, or one generated by another model.
@@
To gauge self-recognition ability we employ two different paradigms, hereafter referred to as Paired and Individual presentation. In Paired presentation, as described in Panickssery et al. (2024), the model is shown two summaries or continuations, labeled 1 and 2, one of which has been generated by the model being tested and the other of which has been generated by a human or another model, and the model is asked to output the label of the summary it thinks it wrote. We include the article or base text from which the summaries or continuations have been generated, along with the instructions given to the writer, as context. The model is shown each pair of texts twice, in different passes, once with its own text first and once with its text second. We capture the log probability assigned to the “1” and “2” tokens. Our accuracy metric simply counts the percentage of times the model gives the self-written text a higher probability that the other-written text across the 2*dataset size presentations.
In Individual presentation, which utilized the same datasets, the model is shown a text and asked to indicate whether it thinks it wrote it, considering the context in which it was generated (again including the instructions and base texts). We captured the probabilities assigned to the “Yes” and “No” tokens, and as our primary metric of accuracy, compute the percentage of times the probability of the “Yes” token after the self-generated text is higher than that of the probability of the “Yes” token after the text generated by the other author (human or AI model) that the model is being tested on its ability to distinguish its own outputs from.
@@ 2.2 Creating the “self-recognition” vector
To form the contrast vector, we identified 734 pairs of texts from across the four datasets on which the model had given highly confident and correct judgments for both self and other texts in the Individual presentation paradigm. We then devised a prompt (see Appendix A.2) that would allow us to solicit a two-alternative-forced-choice judgment from the model while controlling for order and varying the wording of the answer options. We balanced assignment of self/other answers to “A” and “B”, which are always the options and always in that order (the model can get confused otherwise). The model has  68% accuracy, reasonably balanced between self and other; only pairs where it got both the self and other judgment correct were used to form the contrastive prompts to identify the steering vector (n = 290). We captured the activations to the final 10 tokens (starting with the newline before “Answer”, and including the response token) in the residual stream, and took the mean difference between activations to self- and other-written texts, which we normalized for steering.
To remove potential surface-level confounds we formed a “nuisance” vector: we created a series of simple prompts that asked the model to output “A” or “B”, “Yes” or “No”, and a number of variations on “I”/“Me”/“My” and “He”/“She”/“Someone”, and again subtracted and normed the difference between pairs. We then, for every layer, subtracted from the “self-recognition” vector its projection onto the nuisance vector at the corresponding layer.
In addition to test set derived from the datasets described above, we employ a novel test set based on a Quora dataset of question and answer pairs (QA; (Datasets, 2021)). We randomly select pairs (n = 1133) from that dataset where the answers are of moderate length (1K-2K characters) and do not contain self-referential words (“I”, etc, which we expected would be disproportionately subjective and thus ones the model would give stereotyped output to), and neither the question nor the answer contained profanity (to avoid refusals). Then we had Llama3-8b-Instruct generate answers to the chosen questions, and the model and human answers were used as “self” and “other” written texts, respectively.
@@ 3.1.1 Paired presentation paradigm using Llama3-8b-Instruct
Our first experiment tested whether Llama3-8b-Instruct could achieve above-chance accuracy at self recognition in the Paired paradigm across a range of datasets. As shown in Figure 1(a), the model can successfully distinguish its own output from that of humans in all four datasets. In all but the SAD dataset it also discriminates its own outputs from that of other AI models. That the model can do so is expected based on prior positive findings, but the level of success is surprisingly high.
In investigating potential causes of this, we noticed that in all datasets except DOLLY, Llama3-8b-Instruct’s output is longer than that of the human output - and frequently than other model output, and that the model used that information (see Figure 8 in Appendix A.6). We include this as an example of how uncontrolled surface differences may allow the model to use information beyond self-knowledge to pass self-recognition tasks. We then took steps to normalize the lengths of all texts, experimenting with filtering down to generated texts that happened to be the same length across authors, and with trimming all texts to a set length - being careful to trim all texts so as to remove the confound of comparing truncated vs non-truncated texts. Both approaches yielded qualitatively similar results; as the latter afforded us more texts to work with, it is what is used in the results that follow.
@@ 3.1.2 Perplexity
Another way a model might succeed at the task without actual knowledge of its own writing style is to use the text’s perplexity according to its own representations. But although this would indeed provide a good signal, as Table 3 in Appendix A.1 shows, Llama3-8b-Instruct does not appear to be using it. Correlations between outputs and perplexity were modest (<=.16) and inconsistent. Compared with human output, model output had substantially lower perplexity in all four datasets, but in only one of them was the correlation even marginally statistically significant, and across all comparisons that relationship was usually in the wrong direction. In most cases the model did a worse job of judging its own outputs than it would have if it just relied on perplexity.
@@ 3.1.3 Paired presentation paradigm using Llama3-8b-base
If the chat model is using actual knowledge of its own writing style, presumably that exposure would have come in the course of post-training. Thus, that would imply that the base model would fail at this task. And indeed, that was the case, as shown in Figure 2. (Interestingly, the base model actually did worse on the non-length normalized CNN and XSUM datasets than on the length normalized ones; subsequent analysis showed that was at least in part because it was using length as a signal, its own outputs were longer, but it was using longer length as a predictor of other output. Apparently it was picking up on length as something that differentiated texts and simply guessing what it meant, and happening to guess wrong. This implies that the chat model’s use of length might not have been a confound at all, but rather reflected true knowledge of the kind of outputs it generates.
@@ 3.2.1 Isolating a vector
We used contrasting pairs of self- and other-written texts, followed by out-projection of nuisance components (see Methods) to isolate directions in the residual stream that corresponded to the choice of claiming self or other authorship. The final token before the model gave its response was identified as the most promising target. Figure 7 in Appendix A.5 shows the Tuned Lens decoding of it: the positive direction contains a number of tokens related to self-recognition (e.g., “my”, “I”, “match”), while the negative direction contains a number of tokens pointing in the other direction (e.g., “your”, “you”, “other”).
We tested the relevance of this vector for model output by steering (Turner et al., 2023) with it, at a range of multipliers and layers, As can be seen in Figure 3, steering with multipliers in the 3 to 6 range on layers 14-16 was most effective, achieving 100% effectiveness at causing the model to claim authorship, regardless of true authorship, or, when steered in the opposite direction, deny authorship, for both self- and human-written texts.
Figure 3: Detailed steering results for Individual presentation paradigm test set 1.
This pattern was very consistent. Figure 4 combines positive and negative steering for self and other written texts to create a map of aggregate effectiveness for two new datasets, one comprised of test-set texts from the Individual paradigm, and one from the QA dataset. Early layers show little or negative steering effectiveness (high multipliers can lead to degenerate output in the early layers especially), then there’s a sudden, sharp increase in effectiveness at moderate multipliers at layer 14 that peaks at layer 16, and then quickly drops off.
Figure 4: Aggregate steering effectiveness by layer and multiplier in two datasets. 100 = Complete steering effectiveness in the intended direction. Values below 0 are generally degenerate output.
@@ 3.2.4 Causality
Figure 5: Effect of projecting self-recognition vector out of output token on three datasets.
@@ 4 Discussion
Our experiments with zeroing out the vector show that the chat model relies on it to make these self-recognition judgments. Our steering and coloring experiments show that the model’s behavior and perception can be controlled towards or away from self text recognition via manipulation of the vector. The fact that this approach is effective both in a paradigm where the model has to decide whether it wrote a text and in one where it has to decide which text it wrote suggests generality beyond a particular task. We believe this finding holds promise for building resistance to jailbreaks and other model safety risks.
@@ Appendix A Appendix
A.1 Impact of perplexity on self-recognition judgement
Table 3: Perplexity impact, Paired paradigm. S/O Pplx, median ratio of perplexities of self- to other-written texts. %Tot S$<$O Pplx, percentage of pairs in which self-written texts had lower perplexity than other-written texts. PplxRatio-ProbRatio Corr, correlation between self/other perplexity ratios and self/other probability ratios.
@@ A.1 Impact of perplexity on self-recognition judgement
S/O %Tot S$<$O PplX Corr (CI) S/O %Tot S$<$O PplX Corr (CI) S/O %Tot S$<$O PplX Corr (CI) S/O %Tot S$<$O PplX Corr (CI)
@@ A.6 Impact of length on likelihood of self-attribution
Figure 8: Length impact, Paired paradigm. S/O Len, median ratio of the lengths of self- to other-written texts. %Tot $S<O$ Len, percentage of pairs in which self-written texts were shorter than other-written texts. LenRatio-ProbRatio Corr, the correlation between self/other length ratios and self/other averaged probabilities. Odds Ratio is from a logistic regression using length ratio as a predictor.

### v2

@@ Abstract
It has been reported that LLMs can recognize their own writing. As this has potential implications for AI safety, yet is relatively understudied, we investigate the phenomenon, seeking to establish: whether it robustly occurs at the behavioral level, how the observed behavior is achieved, and whether it can be controlled. First, we find that the Llama3-8b–Instruct chat model - but not the base Llama3-8b model - can reliably distinguish its own outputs from those of humans, and present evidence that the chat model is likely using its experience with its own outputs, acquired during post-training, to succeed at the writing recognition task. Second, we identify a vector in the residual stream of the model that is differentially activated when the model makes a correct self-written-text recognition judgment, show that the vector activates in response to information relevant to self-authorship, present evidence that the vector is related to the concept of “self” in the model, and demonstrate that the vector is causally related to the model’s ability to perceive and assert self-authorship. Finally, we show that the vector can be used to control both the model’s behavior and its perception, steering the model to claim or disclaim authorship by applying the vector to the model’s output as it generates it, and steering the model to believe or disbelieve it wrote arbitrary texts by applying the vector to them as the model reads them.
@@ 1 Introduction
Panickssery et al. (2024), utilizing a summary writing/recognition task, report that a number of LLMs, including Llama2-7b-chat, show out-of-the-box (without fine-tuning) self recognition abilities. However, that work focused on the relationship between self-recognition task success and self-preference, rather than the specific means by which the model was succeeding at the task. Laine et al. (2024), as part of a larger effort to provide a foundation for studying situational awareness in LLMs, utilized a more challenging text continuation writing/recognition task and demonstrate self-recognition abilities in several larger models (although not Llama2-7b-chat), but there the focus was on how task success could be elicited with different prompts and in different models. Thus we seek to fill a gap in understanding what exactly models are doing when they succeed at a self recognition task.
We first demonstrate Llama3-8b–Instruct self-recognition task success in a variety of domains. We are particularly interested in distinguishing “true” self recognition of writing - entailing knowledge of one’s own writing style - from “discriminability” - being able to detect consistent differences in the styles of two sets of texts, or being able to identify texts as being more or less “human-like” (i.e., like the pre-training data). To understand whether the model is engaging in “true” self recognition, which would have the implications for AI safety described above, we next attempt to eliminate competing hypotheses.
Having done so, to help understand, and potentially control, the model representations underlying this self-recognition ability, we apply the contrastive pairs method (Turner et al., 2024; Zou et al., 2023) to isolate vectors in the residual stream of the model that are distinctively activated for self- vs human-written texts in the context of a paradigm that prompts the model to make a binary judgment about its authorship of a given text. Via inspection with the Tuned Lens (Belrose et al., 2023) and a series of steering experiments, we then identify one particular vector that appears to be strongly related to the model’s ability to correctly claim or deny authorship. After further work to characterize the information it carries, we then demonstrate that the vector can be applied to the output token to cause the model to assert or deny authorship at essentially 100% rates for new and out-of-distribution texts in this paradigm, and that the model is much less likely to assert authorship when the vector is completely projected out of the residual stream. Having established that the vector directly affects behavior, we then probe whether it can affect perception, by adding it to or subtracting it from the texts being evaluated, and not doing anything to the output token, in both this paradigm and one in which the model is asked to choose which of two texts it wrote. We find a strong effect of this perceptual “coloring” on the model’s behavior.
@@ 2.1 Testing model self-recognition ability
To generate writing samples we employ two different tasks, referred to hereafter as “Summarization” and “Continuation”. In the Summarization task, following Panickssery et al. (2024), models are prompted to generate summaries or other descriptive information about articles, and are subsequently tested on their ability to distinguish those from responses generated by humans or other models. In the Continuation task, following Laine et al. (2024) and using a portion of the dataset they constructed (“SAD”), we show the models the beginning of a piece of human-generated text, then prompt them to continue it in a plausible manner, then probe their ability to distinguish what they generated from the true continuation, or one generated by another model.
@@
To gauge self-recognition ability we employ two different paradigms, hereafter referred to as Paired and Individual presentation. In Paired presentation, as described in Panickssery et al. (2024), the model is shown two summaries or continuations, labeled 1 and 2, one of which has been generated by the model being tested and the other of which has been generated by a human or another model, and the model is asked to output the label of the summary it thinks it wrote. We include the article or base text from which the summaries or continuations have been generated, along with the instructions given to the writer, as context (see Appendix A.11 for further discussion). The model is shown each pair of texts twice, in different passes, once with its own text first and once with its text second. We capture the log probability assigned to the “1” and “2” tokens. Our accuracy metric simply counts the percentage of times the model gives the self-written text a higher probability that the other-written text across the 2*dataset size presentations.
In Individual presentation, which utilized the same datasets, the model is shown a text and asked to indicate whether it thinks it wrote it, considering the context in which it was generated (again including the instructions and base texts). We captured the probabilities assigned to the “Yes” and “No” tokens, and as our primary metric of accuracy, compute the percentage of times the probability of the “Yes” token after the self-generated text is higher than that of the probability of the “Yes” token after the text generated by the other author (human or LLM) that the model is being tested on its ability to distinguish its own outputs from.
@@ 2.2 Creating the “self-recognition” vector
[…]
To remove any potential remaining surface-level confounds relating to output, we formed a “nuisance” vector: we created a series of simple prompts that asked the model to output “A” or “B”, “Yes” or “No”, and a number of variations on “I”/“Me”/“My” and “He”/“She”/“Someone”, ran them through the model and captured activations to the final token before output, and subtracted and normed the activation difference between pairs. We then, for every layer and position, subtracted from the “self-recognition” vector its projection onto the nuisance vector at the corresponding layer.
In addition to the test set derived from the datasets described above, we employ a novel test set based on a Quora dataset of question and answer pairs (QA; (Datasets, 2021)). We randomly select pairs (n = 1133) from that dataset where the answers are of moderate length (1K-2K characters) and do not contain self-referential words (“I”, etc, which we expected would be disproportionately subjective and thus ones the model would give stereotyped output to), and neither the question nor the answer contained profanity (to avoid refusals). Then we had Llama3-8b-Instruct generate answers to the chosen questions, and the model and human answers were used as “self” and “other” written texts, respectively.
@@ 3.1.1 Paired presentation paradigm using Llama3-8b-Instruct
Our first experiment tested whether Llama3-8b-Instruct could achieve above-chance accuracy at self recognition in the Paired paradigm across a range of datasets. As shown in Figure 1(a), the model can successfully distinguish its own output from that of humans in all four datasets. In all but the SAD dataset (bottom right panel) it also discriminates its own outputs from that of other AI models. That the model can do so is expected based on prior positive findings, but the level of success is surprisingly high.
In investigating potential causes of this, we noticed that in all datasets except DOLLY, Llama3-8b-Instruct’s output is longer than that of the human output - and frequently than other model output, and that the model used that information (see Figure 8 in Appendix A.6). We include this as an example of how uncontrolled surface differences may allow the model to use information beyond self-knowledge to pass self-recognition tasks. We then took steps to normalize the lengths of all texts, experimenting with filtering down to generated texts that happened to be the same length across authors, and with trimming all texts to a set character length - being careful to trim all texts so as to remove the confound of comparing truncated vs non-truncated texts. Both approaches yielded qualitatively similar results; as the latter afforded us more texts to work with, it is what is used in the results that follow.
@@ 3.1.2 Perplexity
Another way a model might succeed at the task without actual knowledge of its own writing style is to use the text’s perplexity according to its own representations. But although this would indeed provide a good signal, as Table 3 in Appendix A.1 shows, Llama3-8b-Instruct does not appear to be using it. Correlations between outputs and perplexity were modest (<=0.16) and inconsistent. Compared with human output, model output had substantially lower perplexity in all four datasets, but in only one of them was the correlation even marginally statistically significant, and across all comparisons that relationship was usually in the wrong direction. In most cases the model did a worse job of judging its own outputs than it would have if it just relied on perplexity.
@@ 3.1.3 Paired presentation paradigm using Llama3-8b-base
If the chat model is using actual knowledge of its own writing style, presumably that exposure would have come in the course of post-training. Thus, that would imply that the base model would fail at this task. And indeed, that was the case, as shown in Figure 2. (Interestingly, the base model actually did worse on the non-length-normalized CNN and XSUM datasets than on the length-normalized ones; subsequent analysis showed that was at least in part because it was using length as a signal, its own outputs were longer, but it was using longer length as a predictor of other output. Apparently it was picking up on length as something that differentiated texts and simply guessing what it meant, and happening to guess wrong. This implies that the chat model’s use of length might not have been a confound at all, but rather reflected true knowledge of the kind of outputs it generates.
@@ 3.2.1 Isolating a vector
We used contrasting pairs of self- and other-written texts, followed by out-projection of nuisance components (see Methods) to isolate directions in the residual stream that corresponded to the choice of claiming self or other authorship. The final token before the model gave its response was identified as the most promising target. Figure 7 in Appendix A.5 shows the Tuned Lens decoding of it: the positive direction contains a number of tokens related to self-recognition (e.g., “my”, “I”, “match”), while the negative direction contains a number of tokens pointing in the other direction (e.g., “their”, “different”, “other”).
We tested the relevance of this vector for model output by steering (Turner et al., 2024) with it (adding scaled versions of it to activations in the residual stream during generation), at a range of multipliers and layers, As can be seen in Figure 3, steering with multipliers in the 3 to 6 range on layers 14-16 was most effective, achieving 100% effectiveness at causing the model to claim authorship, regardless of true authorship, or, when steered in the opposite direction, deny authorship, for both self- and human-written texts.
[…]
This pattern was very consistent. Figure 4 combines positive and negative steering for self and other written texts to create a map of aggregate effectiveness for two new datasets, one comprised of test-set texts from the Individual paradigm, and one from the QA dataset. Early layers show little or negative steering effectiveness (high multipliers can lead to degenerate output in the early layers especially), then there’s a sudden increase in effectiveness at moderate multipliers at layer 14 that peaks at layer 16, and then quickly drops off.
Figure 4: Aggregate steering effectiveness by layer and multiplier in two different datasets (left and right). 100 = Complete steering effectiveness in the intended direction. Values below 0 mean that the steered model was less likely to claim (for positive steering) or deny (for negative steering) authorship than the unsteered model, and are generally indicative of degenerate output.
@@ 3.2.4 Causality
[…]
@@ 3.2.5 Coloring
In summary, by contrasting residual stream activations to pairs of model- and human-written texts in a carefully constructed paradigm, we have created a vector that 1) carries information related to asserting self recognition, 2) can be used to steer the model to claim or deny authorship, 3) cannot be used to steer the model towards generic agreement, 4) can be used to steer the model towards more “self”-related outputs during open-ended generation, 5) is preferentially activated when the model reads texts bearing the characteristics of RLHF’d model output, 6) accumulates self-text-recognition-relevant information (in both the chat and base models) and converts it to a decision (in the chat model but not the base model) as it proceeds through layers, 7) is causally related to the model’s ability to assert self-authorship, and 8) can be applied to input tokens to make the model believe they are its own.
@@ 4 Discussion
Our experiments with zeroing out the vector show that the chat model relies on it to make these self-recognition judgments. Our steering and coloring experiments show that the model’s behavior and perception can be controlled towards or away from self-written text recognition via manipulation of the vector. The fact that this approach is effective both in a paradigm where the model has to decide whether it wrote a text and in one where it has to decide which text it wrote suggests generality beyond a particular task. While we here focus on Llama3, we expect that our findings regarding self-recognition ability apply to larger models and ones outside the Llama family. Prior work (Panickssery et al., 2024; Laine et al., 2024) using similar paradigms suggests that RLHF’d Claude and GPT models show comparable self-text recognition abilities to Llama ones, and that abilities increase with scale, and our own experiments with Sonnet 3.5 using our current paradigm (Figure 14) indicate that it has superior self-text recognition abilities to Llama3-8b. We believe that our findings regarding the self-text recognition vector hold promise for building resistance to jailbreaks and other model safety risks. For example, we hypothesize that the vector can be used to prevent users from injecting realistic fake previous responses, as in many-shot jailbreaking (Anil et al., 2024), by adding it to the model’s output tokens and subtracting it from input tokens. It may also be possible to employ it as a sort of warning system: observing endogenous vector activation to arbitrary text can be an alert that the model knows (or does not know) that it or a human is speaking.
@@ Appendix A Appendix
A.1 Impact of perplexity on authorship judgment
Table 3: Perplexity impact, Paired paradigm. S/O, median ratio of perplexities of self- to other-written texts. %Tot S<<<O, percentage of pairs in which self-written texts had lower perplexity than other-written texts. Pplx Corr, correlation between self/other perplexity ratios and self/other probability ratios. Although the perplexity of self-written text was much lower than that of human- and non-Llama-model-written text across datasets, that relationship bore little relationship to the probability the model assigned to having written the texts.
@@ A.1 Impact of perplexity on authorship judgment
S/O %Tot S<<<O PplX Corr (CI) S/O %Tot S<<<O PplX Corr (CI) S/O %Tot S<<<O PplX Corr (CI) S/O %Tot S<<<O PplX Corr (CI)
@@ A.6 Impact of length on likelihood of self-attribution
Figure 8: Length impact, Paired paradigm. S/O Len, median ratio of the lengths of self- to other-written texts. %Tot S<<<O Len, percentage of pairs in which self-written texts were shorter than other-written texts. LenRatio-ProbRatio Corr, the correlation between self/other length ratios and self/other averaged probabilities. Odds Ratio is from a logistic regression using length ratio as a predictor.